# A Proxy Attack-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v2

@@ IV-A Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
We assert that the primary reason for the efficiency of backdoor injection is due to the similarity between benign and corresponding poisoning samples. Typically, poisoning samples exhibit different labels than their clean counterparts, signifying that these pairs share similar attributes yet are categorized differently in dirty-label backdoor attacks. In this context, samples exhibiting a high degree of similarity can be considered challenging samples in the context of poisoning tasks. Consequently, they tend to be more efficient than low-similarity samples for the purpose of backdoor attacks. To substantiate this assumption, we conducted empirical investigations utilizing the CIFAR-10 dataset while adhering to the same experimental settings as detailed in Sec. III-A. The left segment of Figure 5 visually captures the distribution of similarities among 500 randomly chosen samples and 500 samples searched using the FUS method [50]. The result demonstrates that the FUS-searched efficient sample exhibits higher similarity compared to randomly selected samples.
@@ IV-B Threat Model
Algorithm 1 Proxy attack-Free Selection Strategy (PFS)
@@
IV-C Proxy Attack-free Selection Strategy
Drawing inspiration from our analysis of forensic features contributing to efficient data in poisoning-based backdoor attacks, as showcased in Sec. IV-A, we introduce a simple yet efficient sample selection strategy termed the Proxy attack-Free Strategy (PFS) for improving poisoning efficiency. Unlike prior approaches that rely on a proxy attack task for sample selection, our technique capitalizes on the similarity between benign and corresponding poisoning samples, which is only related to the Data (D) and Trigger (T), liberating it from dependence on the settings of actual attack within victim phase. Specifically, we adopt a pre-trained feature extractor $E(\cdot)$ to compute the cosine similarity (cos($\cdot$)) within the feature space between benign and corresponding poisoning samples. Subsequently, we identify the top $m\times r$ most similar samples and randomly select $r$ from this subset as the poisoning set, where $m$ is a hyper-parameter that modulates the diversity of the poisoning samples. The procedure of our methodology is presented in Algorithm 1. Acknowledging that our PFS relies on a feature extractor pre-trained on a clean dataset to compute similarity, we contend that the essence of FUS and PFS diverges significantly. Our PFS identify the forensic features of efficient samples in poisoning-based backdoor attacks, and adopt the similarity and diversity to identify efficient samples, which is disentangled from the actual attack within victim phase. Therefore, the effectiveness of PFS primarily depends on the accuracy of similarity measurement.
@@ IV-C Proxy Attack-free Selection Strategy
IV-D Theoretical Analyses
Additionally, we provide theoretical analyses of the proposed PFS in terms of both active learning and neural tangent kernel.
Perspective of Active Learning. The process of selecting efficient poisoning samples in the context of backdoor attacks can be analogized to the principles of active learning. Active learning (AL) is a machine learning paradigm that efficiently acquires annotated data from an extensive pool of unlabeled data [35]. Its objective is to prioritize human labeling efforts on the most informative data points, leading to improved model performance while reducing data annotation costs. Similarly, the task of selecting efficient poisoning samples in a backdoor attack aims to identify the most informative data points that can optimize the efficiency of the poisoning process, thereby minimizing the costs for the attacker. A pivotal aspect of active learning encompasses the concepts of uncertainty and diversity in the dataset [6]. Methods that rely on uncertainty utilize the predictive confidence of model to select challenging examples, while diversity-based sampling leverages the dissimilarity among data points, often involving clustering techniques. Notably, the combination of uncertainty and diversity strategies has demonstrated significant success in active learning [56, 42]. Similarly, our proposed PFS takes into account both uncertainty (Line 8 in Algorithm 1) and diversity (Line 9 in Algorithm 1) of data points.
Perspective of Neural Tangent Kernel (NTK). To further explain this interesting phenomenon (i.e., similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set emerge as two pivotal factors influencing the efficiency of data in backdoor attacks), we exploit recent studies on NTK [22, 15, 16, 30] for analyzing the sample efficiency of backdoor attacks:
@@ IV-D Theoretical Analyses
In general, Theorem 1 suggests that selecting samples characterized by high similarity between clean ($x_{i}$) and corresponding poisoning samples ($x^{\prime}_{i}$), as well as high similarity between poisoning samples ($x^{\prime}_{i}$) and poisoned testing samples ($x^{\prime}_{t}$), enhances the confidence in predicting poisoned samples to the target class. In the context of poisoning-based backdoor attacks, where training samples share the same distribution as testing samples, high similarity between poisoning samples and poisoned testing samples indicates congruence in the distribution between the poisoning set and the training set, signifying a high level of diversity within the poisoning set. The proof of this theorem can be found in the Sec. A of the Supplementary Materials.
V Experiments
We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec. V-B), Tiny-ImageNet (Sec. V-C), and CIFAR-100 (Sec. E of the Supplementary Materials). We also provide experimental settings in Sec. V-A, comparison with recent sample selection methods in Sec. V-D, time consumption analysis in Sec. V-F, ablation study in Sec. V-G, and experiment against defense method in Sec. V-H.
@@ V Experiments
TABLE I: The attack success rate under different data transforms on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets, Blended, and Optimized are $1\%$, $1\%$, and $0.5\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@ V-A Experimental Settings
We conduct comprehensive experiments encompassing three trigger types§§§Additionally, we consider the recent backdoor attacks such as WaNet [38] and ISSBA [27] in Sec. V-D. (BadNets [14], Blended [4], Optimized [61]), six data transformations (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three different DNN architectures (VGG16 [43], ResNet18 [19], PreAct-ResNet18 [20]). Furthermore, we explore the effects of four optimizers: SGD-0.01SGD-0.01 implies the adoption of the SGD optimizer with an initial learning rate of 0.01., SGD-0.02, Adam-0.001, and Adam-0.002. The employed feature extractor $E(\cdot)$ is ResNet18, pre-trained on the specific dataset using the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, along with two standard data augmentations: random crop and random horizontal flip. Each training iteration spans 70 epochs, conducted with a batch size of 256. In addition, Sec. V-E presents experiments involving various pre-trained feature extractors on the CIFAR-10 dataset. Our findings reveal that even utilizing a commonly used ImageNet pre-trained feature extractor yields notable enhancements compared to the baseline.
We perform a comprehensive comparative analysis, evaluating our method against the widely employed random selection (Random) [14, 4] and FUS [50] approaches. In the FUS search procedure, we utilize the VGG16 architecture, coupled with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer to establish the proxy attack. Conversely, for our proposed method, we utilize pre-trained ResNet models for similarity measurement and set the diversity hyperparameter $m$ to 10 across all three datasets. All other hyperparameters remain consistent with those outlined in Sec. III-A. Additionally, Sec. V-D shows the comparison with recent sample selection methods (FUS [50], OPS [17], Gradient[10], RD Score[49]) on the Tiny-ImageNet dataset.

### v3

@@ III-B Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
We assert that the primary reason for the efficiency of backdoor injection is due to the similarity between benign and corresponding poisoning samples. Typically, poisoning samples exhibit different labels than their clean counterparts, signifying that these pairs share similar attributes yet are categorized differently in dirty-label backdoor attacks. In this context, samples exhibiting a high degree of similarity can be considered challenging samples in the context of poisoning tasks. Consequently, they tend to be more efficient than low-similarity samples for the purpose of backdoor attacks. To substantiate this assumption, we conducted empirical investigations utilizing the CIFAR-10 dataset while adhering to the same experimental settings as detailed in Sec. IV-A. The left segment of Figure 2 visually captures the distribution of similarities among 500 randomly chosen samples and 500 samples searched using the FUS method [63]. The result demonstrates that the FUS-searched efficient sample exhibits higher similarity compared to randomly selected samples.
@@ III-C Threat Model
Algorithm 1 Proxy attack-Free Strategy (PFS)
@@
III-D Proxy attack-Free Strategy
Drawing inspiration from our analysis of forensic features contributing to efficient data in backdoor attacks, as showcased in Sec. III-B, we introduce a simple yet efficient sample selection strategy termed the PFS to improve poisoning efficiency. Unlike prior approaches that rely on a proxy attack task for sample selection, our technique leverages the similarity between benign and corresponding poisoned samples, which depends solely on the Data (D) and Trigger (T), freeing it from dependence on the settings of the actual attack during the victim phase. We adopt a pre-trained feature extractor $E(\cdot)$ to compute the cosine similarity $cos(\cdot)$ within the feature space between benign and corresponding poisoning samples. Specifically, for a given benign image $x_{i}\in\mathcal{D}$, Backdoor trigger $t$, and Trigger function $\mathcal{T}$, the corresponding poisoned image is defined as $x^{\prime}_{i}=\mathcal{T}(x_{i},t)$. Consequently, the similarity between benign and corresponding poisoned samples in the feature space can be computed by:
$s_{i}=cos(E(x_{i}),E(x^{\prime}_{i}))\end{aligned}\text{.}$ (2)
The similarity set is denoted as $S=\{s_{1},\cdots,s_{i},\cdots,s_{N}\}$, where $N$ is the size of the benign training set, which contains the top $m\times r$ most similar samples from the benign training set $\mathcal{D}$ based on their similarity scores. This step filters out inefficient samples. Then, to ensure diversity, we randomly select $r$ samples from the coarse poisoning set $\mathcal{\hat{P}}$ to form the final poisoning set, where $m$ is a hyper-parameter that modulates the diversity of the poisoning samples.
In summary, the procedure of our methodology is presented in Algorithm 1. While PFS relies on a feature extractor pre-trained on a clean dataset to compute similarity, we argue that the essence of FUS and PFS diverges significantly. PFS identifies the forensic features of efficient samples in poisoning-based backdoor attacks and uses similarity and diversity to select efficient samples, independent of the actual attack within the victim phase. Therefore, the effectiveness of PFS primarily depends on the accuracy of similarity measurement.
@@ III-D Proxy attack-Free Strategy
III-E Theoretical Analyses
Additionally, we provide theoretical analyses of the proposed PFS in terms of both active learning (Sec. A of the Supplementary Materials.) and neural tangent kernel.
Perspective of Neural Tangent Kernel (NTK). To explain the interesting phenomenon (i.e., similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set emerge as two pivotal factors influencing the efficiency of data in backdoor attacks), we exploit recent studies on NTK [27, 18, 19, 36] for analyzing the sample efficiency of backdoor attacks:
@@ III-E Theoretical Analyses
Theorem 1 indicates that the confidence in predicting poisoned samples to the target class, $\mathbb{E}_{{x^{\prime}_{t}}}\left[f_{\theta}({x^{\prime}_{t}})\right]$ is inversely proportional to the values of ${x}^{\prime}_{{i}}-{x}^{\prime}_{{t}}$ and ${x}^{\prime}_{{i}}-{x}_{{i}}$. This suggests that selecting samples characterized by high similarity between benign ($x_{i}$) and corresponding poisoning samples ($x^{\prime}_{i}$), as well as high similarity between poisoning samples ($x^{\prime}_{i}$) and poisoned testing samples ($x^{\prime}_{t}$), enhances the confidence in predicting poisoned samples to the target class. In the context of poisoning-based backdoor attacks, where training samples share the same distribution as testing samples, high similarity between poisoning samples and poisoned testing samples indicates congruence in the distribution between the poisoning set and the training set, signifying a high level of diversity within the poisoning set. Therefore, this theorem indicates that the similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set are pivotal factors influencing the efficiency of data in backdoor attacks. The proof of this theorem can be found in the Sec. B of the Supplementary Materials.
IV Degradation of proxy attack-based Selection
This section demonstrates that proxy attack-based sample selection methods experience significant degradation in sample efficiency when there is a substantial difference between the settings used for proxy poisoning attacks during the attacker phase and those employed in the actual poisoning process during the victim phase. The purpose of this section is to highlight the limitations of sample selection methods that rely on proxy attacks, thereby motivating the design of a proxy attack-free sample selection strategy for backdoor attacks. To better understand our key observations, we will first review the sample selection methods based on proxy attacks.
Proxy Attacks-based Sample Selection. Recent investigations [63, 69, 13] have highlighted that diverse samples contribute differently to the backdoor injection and a particular subset with efficient data has higher attack success rate. Notably, [63, 13] claim that the efficacy of poisoning samples is intertwined with instances of ”forgetting events” during the backdoor injection process. Although FUS [63], Gradient [13], and OPS [21] have acquired remarkable performance, they use a proxy attack processing to find the efficient subset.
As illustrated in Figure 1, poisoning-based backdoor attacks contain two phases: Attacker combines data (D) and trigger (T) to build poisoned data (D+T) and release it. Victims download the released data and adopt it to train infected DNNs. Notably, during the training process, various data transformations/augmentations (F) and hyperparameters (H)***In this context, H encompasses various training settings, encompassing elements like the model architecture, optimizer choice, and training hyperparameters. have been applied by victims. Generally, attacker has no access to any information of the victim phase and only performs the poisoning injection at the attack phase. However, proxy attack-based sample selection methods need to select samples with the help of the complete attack process (both attacker and victim phases). We argue that all parts involved in backdoor attacks are the factors that affect the efficiency of poisoning samples, containing D, T, F, and H. When there contains some gaps (e.g. Transform (F) and Hyperparameters (H) in Figure 1) between proxy poisoning attack and actual poisoning process of victims, the poisoning effectiveness of the selected samples may be decreased due to the shortcut nature of backdoor learning. Following, we will explore what will happen if the actual poisoning process is different from the proxy poisoning attack in proxy attack-based sample selection methods.
IV-A Experimental Settings
To investigate the proposed question within this section, we engage in a series of experiments utilizing the VGG16 model [55] as the victim model. These experiments are conducted on the CIFAR-10 and Tiny-ImageNet datasets. Our chosen config encompasses the SGD optimizer, initially set with a learning rate of 0.01, a momentum value of 0.9, and a weight decay coefficient of 5e-4. Notably, we incorporate two widely adopted data transformations—random cropping and random horizontal flipping. The training is performed for 70 epochs with a batch size of 256. Poisoning samples $x^{\prime}_{i}$ are generated using a blending strategy [6] in which the clean samples $x_{i}$ are blended with a trigger image $t$ using the equation $x^{\prime}_{i}=\mathcal{T}(x_{i},t)={{\lambda}{\cdot}{t}}+(1-{\lambda}){\cdot}{x_{i}}$, where ${\lambda}$ is set to 0.15. We set category 0 as the target for the attack, and we set the poisoning rate with $r=0.01$—equating to a poisoning set size of $P=500$. The process involves 15 iterations of selection, aligning with the settings of [63]. In accordance with previous literature [32], the attack success rate is defined as the probability of classifying a poisoned test data to the target label. Unless stated otherwise, all experiments in this section follow the aforementioned settings. To explore the degeneration in poisoning efficiency within selected samples due to disparities between the proxy poisoning attack and the actual poisoning process of victims, we compare the Filtering-and-Updating Strategy (FUS) with SELF-FUS. SELF-FUS represents an ideal scenario for FUS, assuming the attacker possesses complete information about both the attacker and victim phases during poisoning-based backdoor attacks—a scenario that doesn’t align with the threat model in this study. Therefore, proxy poisoning attack of attacker phase is consistent with the actual poisoning process of victim phase in SELF-FUS, which is not guaranteed in the FUSSimilarly, for Gradient [13] and OPS [21], we compare the performance of these methods with their respective idealized versions, SELF-Gradient and SELF-OPS. SELF-Gradient and SELF-OPS represent scenarios where the attacker has complete information about both the attack and victim phases during poisoning-based backdoor attacks..
IV-B Empirical Studies
Impact of Different Data Transformations (F) on Proxy Attack-based Sample Selection Efficacy. We investigated how F affect the effectiveness of proxy attack-based sample selection methods. To conduct our analysis, we used the experimental settings described previously and followed them exactly. Specifically, we configured the data transformations for the proxy poisoning attack during the sample selection phase (FUS) as ”RandomCrop” and ”RandomHorizontalFlip”. Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of transformations, including ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter”. Our results, depicted in Figure 3, indicate that different data transforms (F) decrease the effectiveness of proxy attack-based sample selection methods. An exception to this pattern emerges in cases where the settings for both the proxy and actual attacks align closely, exemplified by the ”1+2” searching and ”1” attacking scenario. This particular outcome is logical and can be attributed to the inherent margin of error.
Figure 4: The ASR in scenarios where the hyperparameters (H) used during the proxy poisoning attack within attacker phase in FUS is different from the actual poisoning process within victim phase. The coordinates ”V-13, SGD, 0.01”IV-B on the horizontal axis symbolize the utilization of VGG-13 as the architecture and SGD with an initial learning rate of 0.1 as the optimizer. Each reported result is an average computed from five separate runs.
Figure 5: The attack success rate in situations where both the data transformation (F) and the hyperparameter (H) used during the actual attack within the victim phase is different from that used within the search process with Gradient [13], OPS [21], and FUS [63] on Tiny-ImageNet datset. The horizontal coordinates are labeled as ”0”, ”1”, ”2”, ”3”, and ”4”, representing different combinations of data transforms and hyperparameters. All results are computed as the mean of five different runs.
Effect of Different Hyperparameters (H) on Proxy Attack-based Sample Selection Efficiency. In this section, We thoroughly investigated the impact of different H on the effective data efficiency searched by FUS [63]. Adhering to the aforementioned experimental framework, we configured the hyperparameter combination for the proxy poisoning attack during the sample selection phase as ”V-16, SGD, 0.01”. Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of H, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”In this context, V-13, V-16, R-18, and P-18 denote VGG13, VGG16, ResNet18, and PreActResNet18, respectively. . Figure 4 indicates that different H decrease the effectiveness of proxy attack-based sample selection methods. Once again, we observed anomalous occurrences when the parameters for both proxy searching and actual attacks closely converged, as illustrated by the scenario involving ”V-16-SGD-0.01” proxy searching and subsequent ”V-13-SGD-0.01” attacks within the victim phase.
Summary Our findings underscore the susceptibility of proxy attack-based sample selection techniques to shifts in data transforms and hyperparameters deployed during the victim phase. Notably, SELF-FUS consistently outperforms FUS across various scenarios, with the exception of outliers observed in both Figure 3 and Figure 4. These outliers, marked by the ”1+2” searching and ”1” attacking in Figure 3, as well as the ”V-16-SGD-0.01” searching and ”V-13-SGD-0.01” attacking in Figure 4, share closely aligned proxy and actual attack settings. To further validate our hypothesis, we continue to simulate more realistic attack scenarios that introduce disparities in both data transformations (F) and hyperparameters (H) between proxy and actual attacks. Figure 5 presents the ASR comparison between FUS and SELF-FUS, Gradient and SELF-Gradient, and OPS and SELF-OPS on the Tiny-ImageNet dataset. The notable differences in ASR highlight the significant impact of proxy attack-based methods on sample efficiency.
@@ IV-B Empirical Studies
TABLE I: The attack success rate under different optimizers on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets, Blended, ISSBA, and Optimized are $1\%$, $1\%$, $1\%$, and $0.5\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
Trigger Optimizer Model
VGG16 ResNet18 PreAct-ResNet18
Random PFS FUS FUS+PFS Random PFS FUS FUS+PFS Random PFS FUS FUS+PFS
ISSBA SGD-0.01 0.741 0.801 0.828 0.856 0.769 0.862 0.847 0.882 0.779 0.851 0.838 0.868
SGD-0.02 0.763 0.839 0.834 0.857 0.806 0.874 0.870 0.889 0.803 0.880 0.878 0.894
Adam-0.0005 0.795 0.826 0.819 0.838 0.822 0.891 0.885 0.902 0.820 0.886 0.872 0.899
Adam-0.001 0.811 0.862 0.858 0.869 0.827 0.883 0.875 0.901 0.822 0.879 0.861 0.902
Avg 0.778 0.832 0.835 0.855 0.806 0.878 0.869 0.894 0.806 0.874 0.862 0.891
TABLE II: The attack success rate under different data transforms on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets, Blended, ISSBA, and Optimized are $1\%$, $1\%$, $1\%$, and $0.5\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@
ISSBA None 0.712 0.782 0.779 0.795 0.729 0.841 0.801 0.852 0.779 0.862 0.833 0.882
RandomCrop 0.708 0.779 0.776 0.802 0.733 0.811 0.809 0.832 0.744 0.824 0.819 0.852
RandomHorizontalFlip 0.732 0.834 0.841 0.860 0.792 0.883 0.852 0.885 0.791 0.881 0.849 0.899
RandomRotation 0.709 0.766 0.772 0.791 0.753 0.819 0.812 0.847 0.754 0.846 0.828 0.852
ColorJitter 0.681 0.763 0.731 0.792 0.689 0.772 0.739 0.838 0.681 0.769 0.750 0.877
RandomCrop+ RandomHorizontalFlip 0.750 0.802 0.834 0.858 0.781 0.857 0.844 0.882 0.783 0.842 0.839 0.869
Avg 0.715 0.788 0.789 0.816 0.746 0.831 0.810 0.856 0.755 0.837 0.820 0.872
@@
V Experiments
We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec. V-B), Tiny-ImageNet (Sec. V-C), and CIFAR-100 (Sec. F of the Supplementary Materials). We also provide experimental settings in Sec. V-A, comparison with recent sample selection methods in Sec. V-D, time consumption analysis in Sec. V-F, ablation study in Sec. V-G, and experiment against defense method in Sec. V-H.
@@ V-A Experimental Settings
We conduct comprehensive experiments encompassing four triggers§§§Additionally, we consider the recent backdoor attacks such as WaNet [50], ISSBA [33], and BATT [67] in Sec. V-D. (BadNets [17], Blended [6], Optimized [76], and ISSBA [33]), six data transformations (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three different DNN architectures (VGG16 [55], ResNet18 [23], PreAct-ResNet18 [24]). Furthermore, we explore the effects of four optimizers: SGD-0.01SGD-0.01 implies the adoption of the SGD optimizer with an initial learning rate of 0.01., SGD-0.02, Adam-0.001, and Adam-0.002. The employed feature extractor $E(\cdot)$ is ResNet18, pre-trained on the specific dataset using the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, along with two standard data augmentations: random crop and random horizontal flip. Each training iteration spans 70 epochs, conducted with a batch size of 256. In addition, Sec. V-E presents experiments involving various pre-trained extractors on the CIFAR-10 dataset. Our findings reveal that even utilizing a commonly used ImageNet pre-trained extractor yields notable enhancements compared to the baseline.
We perform a comprehensive comparative analysis, evaluating our method against the widely employed random selection (Random) [17, 6] and FUS [63] approaches. In the FUS search procedure, we utilize the VGG16 architecture, coupled with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer to establish the proxy attack. Conversely, for our proposed method, we utilize pre-trained ResNet models for similarity measurement and set the diversity hyperparameter $m$ to 10 across all three datasets. All other hyperparameters remain consistent with those outlined in Sec. IV-A. Additionally, Sec. V-D shows the comparison with recent sample selection methods (FUS [63], OPS [21], Gradient[13], RD Score[62]) on the Tiny-ImageNet dataset.